OpenStudy (anonymous):

What is the value of 106 in standard form? 0.00001 0.000001 1,000,000 100,000

OpenStudy (theeric):

I think that number is supposed to be \(10^6\), or 10^6. Do you have any guesses?

OpenStudy (anonymous):

nope can u askk pp

OpenStudy (theeric):

I know, though... \(10^6\) is another way of saying \(10 \times 10 \times 10 \times 10 \times 10 \times 10\). That's what the notation means. And look at \(10\times 10=100\). Then \(100*10=1,000\). You are "tacking on" a zero, for each \(10\) that you multiply. So how many zeroes will you have if you multiply \(6\) of them?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

so its d

OpenStudy (theeric):

Standard form, here, means just the plain old number \(\sf{by\ itself}\).

OpenStudy (anonymous):

so its d

OpenStudy (theeric):

If you're not sure, use a calculator!

OpenStudy (anonymous):

its scientific notation right?

OpenStudy (anonymous):

ye

OpenStudy (theeric):

I guess "standard form" is scientific notation in Britain, but none of the answers are in scientific form.

OpenStudy (anonymous):

can yall hll for a nother ques

OpenStudy (theeric):

Have you settled on an answer, @divalove27 ?

OpenStudy (anonymous):

ye

OpenStudy (theeric):

Which one? Did you use a calculator to be sure?

OpenStudy (anonymous):

yup

OpenStudy (theeric):

Cool! :)

OpenStudy (anonymous):

cann u hllp me with more if u mind

OpenStudy (theeric):

Not D. Sure, just post another question so others can see it as well.

OpenStudy (anonymous):

kk thxx